# Importance sampling for Bayesian inference: polynomial-dimension dependent error bounds

### 02 Apr 2026 v1

@@ Abstract
Many Bayesian inference problems involve high-dimensional models where the performance of standard importance sampling (IS) methods often degrades rapidly as the dimensionality increases. Classical analyses of IS typically rely on the assumption that observations are arbitrary but fixed (i.e., deterministic), thereby neglecting the probabilistic structure that the Bayesian model induces on the data. In this paper, we adopt the perspective that observations are themselves random variables whose distribution is governed by the underlying model. Within this probabilistic framework, we identify a model-dependent function, referred to as the link function, which connects the fixed- and random-observation formulations. We provide a characterization of the $L^{2}$ Monte Carlo estimation error: specifically we show that the $L^{2}$ error bounds are finite and converge at the standard Monte Carlo rate ${\mathcal{O}}(N^{-1/2})$, for arbitrarily large dimension, if and only if the link function is Bochner integrable. This result reveals the fundamental quantity controlling the error of the approximation and establishes a mechanism to manage the dependence on the model state dimension. Consequently, our approach provides a principled way to alleviate the challenges of high dimensionality, offering insights that transcend the worst-case analyses dominant in the existing literature. Finally, we derive explicit analytical examples of the dimensional scaling of the associated errors for several model classes, including linear–Gaussian systems and models with bounded observation functions.
@@ 1.3 Contributions
In this paper, we investigate the role of the dimension of the state (denoted $d_{x}$) in the error bounds achieved by IS methods for Bayesian inference. Unlike much of the previous literature, which assumes fixed observations, we treat the data as r.v.s generated by the model. This approach has been employed in [28] in the context of study of stability for the optimal filter [12]. However, to the best of our knowledge, it has not been applied to IS schemes or Monte Carlo approximations. This natural probabilistic setting allows us to define random posterior measures and, correspondingly, random error bounds, enabling the use of functional-analytic and measure-theoretic tools to exploit the model structure. We first derive $L^{1}$ and $L^{2}$ error bounds for general Bayesian models under random observations and explicitly identify their dependence on the dimension of the model. A central theoretical contribution of the paper is a characterization of the $L^{2}$ Monte Carlo error of the IS estimator in this setting. We show that the canonical $\mathcal{O}(N^{-1/2})$ convergence rate holds if and only if a certain model-dependent function (refereed as the link function) introduced in the analysis is square-integrable with respect to the joint law of the observations and the prior (formalized later using Bochner integrability). This integrability condition emerges as the fundamental quantity governing the error of the approximation. In particular, it provides a direct mechanism to control the dependence of the error bounds on the state dimension $d_{x}$, and in some models even on the observation dimension $d_{y}$, as we explicitly demonstrate in Section 4.1.1. Our analysis relies on the notion of Bochner integrability from functional analysis [39], which, to our best knowledge, has not been previously used for the study of IS estimators.
@@
To derive these results, we first bridge the gap between abstract functional-analytic conditions and practical Bayesian modeling by establishing a concrete, easily verifiable criterion for the integrability of the link function. Specifically, for likelihoods belonging to elliptically symmetric families [15], which encompass diverse noise models such as Gaussian, Laplace, Student-$t$ and Cauchy distributions. We prove that the high-dimensional Bochner integrability condition reduces to the finiteness of a simple, one-dimensional radial integral for models with bounded observations functions.
This reduction allows us to provide explicit analytical constructions showing that, approximation errors grow at most polynomially with the state dimension $d_{x}$. Notably, we identify cases where the polynomial degree is zero, implying error bounds that hold uniformly over $d_{x}$. These findings demonstrate that the exponential complexity typically associated with the curse of dimensionality can be overcome in some practically relevant settings, such as those involving sensors with limited operational ranges.
@@ 1.4 Outline of the Paper
The paper is organized as follows. Section 2 formally describes the probabilistic model and the classical importance sampling algorithm. In Section 3, we analyze the error bounds attained for general models with random observations. The general results obtained in that framework are then particularised to some practically relevant examples in Section 4. Section 5, is devoted to a comparison of the results obtained in Sections 3 and 4 with recent worst case analyses of IS. Finally, in Section 6, we summarise our main findings and discuss possible venues for future research. Technical proofs and some additional results are provided in the appendices. We conclude this introduction with a brief summary of the notation used throughout the paper.
@@ 2.1 Model
We refer to a pair $\mathcal{M}^{d_{x}}=(\pi_{0},g^{d_{x}})$ as a model for Bayesian inference. Throughout this work, we consider a family of models $\{\mathcal{M}^{d_{x}}\}_{d_{x}\in\mathbb{N}}$, indexed by the dimension $d_{x}$ of the state variable $X$. Intuitively, the models in the family $\{\mathcal{M}^{d_{x}}\}_{d_{x}\in\mathbb{N}}$ may be expected to be “of the same type”, i.e., similar in structure and with differences due mainly to the increasing dimension $d_{x}$. An explicit example (of a family of linear and Gaussian models) is given in Section 4.1. Neverheless, the conditions required by the main theorems in Sections 3 and 4 actually allow for families where the individual models can be rather different. The dimension of the observation space, $d_{y}$, is assumed to be fixed unless stated otherwise.
@@ 2.2 Importance sampler
For clarity, we carry out our analysis for a simpler version of this general algorithm where the prior is used as the proposal, i.e. $\nu=\pi_{0}$. The resulting scheme is displayed as Algorithm 2. Hereafter we refer to this specific procedure as standard IS.
@@ 3 Random observations
In this section, we adopt the viewpoint that $Y$ itself is a r.v. generated by the model ${\mathcal{M}}^{d_{x}}$ instead of use it as given (fixed) data. With this perspective, the likelihood function, the normalization constant and the posterior measure also become random objects. In particular, the quantity $\big|\pi_{Y}(f)-\pi_{Y}^{N}(f)\big|$ becomes a $Y$-dependent random variable. The subscript Y emphasizes the randomness of the observation, the same notation is used hereafter for other observation-dependent objects. To derive an upper bound on this random error, we take the expectation with respect to the law of $Y$ in order to obtain an averaged error. This leads us naturally to examine how integration over the random observation is performed. The central tool for this development is the Bochner integrability [14] of a specially constructed function, which acts as a link between the fixed and random observation formulations.
@@ 3.1 Random posterior measure
We aim to derive an upper bound for the approximation error in (5), averaged over the observations, where the expectation is taken with respect to the joint law of the random samples $x^{1},\ldots,x^{N}$ and the observation $Y$. Much of the analysis to be elaborated in this section depends on two objects that we introduce next:
@@
For any test function $\varphi\in L^{2}(\pi_{0})$, the posterior expectation under the random observation $Y$ is given by
$\pi_{Y}(\varphi)=\int_{\mathcal{X}}\varphi(x)\,\pi_{Y}({\sf d}x)=\mathbb{E}\!\left[\varphi(X)\mid\mathcal{G}\right],$ (11)
@@ 3.3 Approximation errors
    $\ell^{\,d_{x}}\in L^{2}(\mathcal{Y};L^{2}(\pi_{0}))$, that is, $\ell^{\,d_{x}}$ is Bochner $2$-integrable.
@@
$N\;\geq\;\left(\frac{P_{n,p}(d_{x})\|f\|_{\infty}}{\varepsilon}\right)^{2}.$ (16)
@@
uniformly with respect to the dimension state $d_{x}$.
@@ 4 Special cases
In this section, we examine specific scenarios under which the assumptions of Theorem 3.3 and Corollary 3.4 are satisfied, that is, when the inequalities (15) and (17) holds. It is worth emphasizing that the bounds derived here are not intended to be tight; rather, our goal is to demonstrate that polynomial and uniform error bounds for the IS estimator $\pi_{Y}^{N}$ can be established as functions of the state dimension $d_{x}$. We focus on the case $p=2$, which subsumes the case $p=1$.
@@
Observe that
$\|\ell_{y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}=\frac{\pi_{0}([g_{y}^{d_{x}}]^{2})}{\pi_{0}(g_{y}^{d_{x}})^{2}},$
which is directly related to the $\chi^{2}$-divergence between the posterior and the prior distributions. This quantity is also related to the relative variance of the Monte Carlo estimator of the normalization constant $Z\mathrel{\mathop{\ordinarycolon}}=\pi_{0}(g_{y}^{d_{x}})$ when sampling from the prior (see Appendix B) and has already been found to be relevant in the analysis of IS methods; see, e.g., [1, 2].
@@
$\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{p}\right]=\int_{\mathcal{Y}}\|\ell_{y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{p}\,\eta(\mathrm{d}y)=\int_{\mathcal{Y}}\frac{\|g_{y}^{d_{x}}\|_{L^{2}(\pi_{0})}^{p}}{\pi_{0}(g_{y}^{d_{x}})^{p-1}}\,{\sf d}y.$ (18)
@@
$\hskip 49.79231pt\int_{\mathcal{Y}}\|\ell_{y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\,\eta(\mathrm{d}y)=\int_{{\mathcal{X}}}\left(\int_{\mathcal{Y}}\frac{g_{y}^{d_{x}}(x)^{2}}{\pi_{0}(g_{y}^{d_{x}})}\,\mathrm{d}y\right)\pi_{0}(\mathrm{d}x).$ (19)
@@
$\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]<\infty.$ (20)
The previous discussion shows that the key condition required for the dimension-dependent error bounds of Theorem 3.3 reduces to the finiteness of the moment in (20). In other words, the error of the IS estimator is governed by the integrability of the link function $\ell_{Y}^{d_{x}}$ as a random element of $L^{2}(\pi_{0})$. In the remainder of this section, we verify this condition for several relevant observation models. The analysis illustrates how the structure of the likelihood determines the dependence of the error bounds on the state dimension $d_{x}$, and shows that polynomial, or even uniform, bounds can arise in common statistical models. We begin with the linear Gaussian model, since in this case all the relevant distributions can be computed analytically, regardless of the dimensionality of the problem.
@@ 4.1 Linear Gaussian model
$Y=A^{d_{x}}X+V,$ (21)
where $A^{d_{x}}\in\mathbb{R}^{d_{y}\times d_{x}}$ is a fixed matrix and $V\sim\mathcal{N}(0,R)$ is a zero-mean Gaussian noise vector with positive definite covariance matrix $R\in\mathbb{R}^{d_{y}\times d_{y}}$, independent of $X$ and also independent of it dimension $d_{x}$. As a result, the r.v. $Y$ is Gaussian with law
$Y\sim\mathcal{N}({\sf d}y;\mu_{y},\Sigma_{y}),$
@@
$\mu_{y}=A^{d_{x}}\mu_{x}^{d_{x}},\qquad\Sigma_{y}=A^{d_{x}}\Sigma_{x}^{\,d_{x}}[A^{d_{x}}]^{\top}+R.$ (22)
@@
We now show that, for the linear Gaussian model described above, there is a finite constant ${\mathcal{K}}_{2}$, independent of the state-space dimension $d_{x}$, that satisfies the inequality (17) (which implies that the assumptions of Corollary 3.4 are satisfied). First, let us note that, since both $A^{d_{x}}\Sigma_{x}^{\,d_{x}}[A^{d_{x}}]^{\top}$ and $R$ are positive definite, it follows that $\Sigma_{y}$ in (22) is positive definite.
In addition, the likelihood function $g_{y}^{d_{x}}(x)$ corresponds to the Gaussian pdf of $Y$ conditional on $X=x$, namely,
@@
$g_{y}^{d_{x}}(x)^{2}=\frac{1}{(2\uppi)^{d_{y}}|R|}\exp\!\left(-\frac{1}{2}(y-A^{d_{x}}x)^{\top}(2R^{-1})(y-A^{d_{x}}x)\right),$
which is proportional to a Gaussian density with covariance $\tfrac{1}{2}R$. Consequently,
$\int_{\mathcal{X}}g_{y}^{d_{x}}(x)^{2}\,\pi_{0}(\mathrm{d}x)=\frac{1}{(2\uppi)^{d_{y}}|R|}\mathcal{N}(y;\mu_{y},\Sigma_{y}-\tfrac{1}{2}R).$ (24)
Using (23) and (24), we realise that in order to satisfy the integrability condition (20) it suffices to find a finite bound for the expectation
$\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]=\frac{1}{(2\uppi)^{d_{y}}|R|}\int_{\mathcal{Y}}\frac{\mathcal{N}(y;\mu_{y},\Sigma_{y}-\tfrac{1}{2}R)}{\mathcal{N}(y;\mu_{y},\Sigma_{y})}\,{\sf d}y.$ (25)
Let $S_{1}\mathrel{\mathop{\ordinarycolon}}=\Sigma_{y}$ and $S_{2}\mathrel{\mathop{\ordinarycolon}}=\Sigma_{y}-\tfrac{1}{2}R$. The integrand in (25) can be rewritten as
$\frac{\mathcal{N}(y;\mu_{y},S_{2})}{\mathcal{N}(y;\mu_{y},S_{1})}=\left(\frac{|S_{1}|}{|S_{2}|}\right)^{1/2}\exp\!\left(-\tfrac{1}{2}(y-\mu_{y})^{\top}(S_{2}^{-1}-S_{1}^{-1})(y-\mu_{y})\right).$
@@
$S_{1}=\Sigma_{y}\succ\Sigma_{y}-\tfrac{1}{2}R=S_{2},$
@@
(hence, $S^{-1}$ is positive definite). The integrand in (25) is thus proportional to the density of a Gaussian distribution with mean $\mu_{y}$ and covariance $S$. As a consequence, the integral in (25) can be computed exactly to yield
$\int_{\mathcal{Y}}\frac{\mathcal{N}(y;\mu_{y},S_{2})}{\mathcal{N}(y;\mu_{y},S_{1})}\,{\sf d}y=\left(\frac{1}{(2\uppi)^{d_{y}}|S|}\frac{|S_{1}|}{|S_{2}|}\right)^{1/2}.$ (27)

### 28 May 2026 v2

@@ Abstract
Many Bayesian inference problems involve high-dimensional models where the performance of standard importance sampling (IS) methods often degrades rapidly as the dimensionality increases. Classical analyzes of IS typically rely on the assumption that observations are arbitrary but fixed (i.e., deterministic), thereby neglecting the probabilistic structure that the Bayesian model induces on the data. In this paper, we adopt the perspective that observations are themselves random variables whose distribution is governed by the underlying model. Within this probabilistic framework, we identify a model-dependent function, referred to as the link function, which connects the fixed- and random-observation formulations. We provide a characterization of the $L^{2}$ Monte Carlo estimation error: specifically we show that the $L^{2}$ error bounds are finite and converge at the standard Monte Carlo rate ${\mathcal{O}}(N^{-1/2})$, for arbitrarily large dimension, if and only if the link function is Bochner integrable. This result reveals the fundamental quantity controlling the error of the approximation and establishes a mechanism to manage the dependence on the model state dimension. Consequently, our approach provides a principled way to alleviate the challenges of high dimensionality, offering insights that transcend the worst-case analyzes dominant in the existing literature. Finally, we derive explicit analytical examples of the dimensional scaling of the associated errors for several model classes, including linear–Gaussian systems and models with bounded observation functions.
@@ 1.3 Contributions
In this paper, we investigate the role of the dimension of the state $X$ (denoted $d_{x}$) in the error bounds achieved by IS methods for Bayesian inference. Unlike much of the previous literature, which assumes fixed observations, we treat the data as r.v.s generated by the model. This approach has been employed in [29] for the study of stability of the optimal filter [14]. However, to the best of our knowledge, it has not been applied to IS schemes or Monte Carlo approximations. This natural probabilistic setting allows us to define random posterior measures and, correspondingly, random error bounds, enabling the use of functional-analytic and measure-theoretic tools to exploit the model structure. We first derive $L^{1}$ and $L^{2}$ error bounds for general Bayesian models under random observations and explicitly identify their dependence on the dimension of the model. A central theoretical contribution of the paper is a characterization of the $L^{2}$ Monte Carlo error of the IS estimator in this setting. We show that the canonical $\mathcal{O}(N^{-1/2})$ convergence rate holds if and only if a certain model-dependent function (refereed as the link function) introduced in the analysis is square-integrable w.r.t. the joint law of the observations and the prior (formalized using Bochner integrability). This integrability condition emerges as the fundamental quantity governing the error of the approximation. In particular, it provides a direct mechanism to control the dependence of the error bounds on the state dimension $d_{x}$. Our analysis relies on the notion of Bochner integrability from functional analysis [40], which, to the best of our knowledge, has not been previously used for the study of IS estimators.
@@
To derive these results, we first bridge the gap between abstract functional-analytic conditions and practical Bayesian modeling by establishing a concrete, easily verifiable criterion for the integrability of the link function. Specifically, for likelihoods belonging to elliptically symmetric families [16] (which encompass diverse noise models such as Gaussian, Laplace, Student-$t$ and Cauchy distributions) we prove that the high-dimensional Bochner integrability condition reduces to the finiteness of a simple, one-dimensional radial integral for models with bounded observations functions.
This reduction allows us to provide explicit analytical constructions showing that approximation errors grow at most polynomially with the state dimension $d_{x}$. Notably, we identify cases where the polynomial degree is zero, implying error bounds that hold uniformly over $d_{x}$. These findings demonstrate that the exponential complexity typically associated with the curse of dimensionality can be mitigated in some practically relevant settings, such as those involving sensors with limited operational ranges.
@@ 1.4 Outline of the Paper
The paper is organized as follows. Section 2 formally describes the probabilistic model and the classical importance sampling algorithm. In Section 3, we analyze the error bounds attained for general models with random observations. The general results obtained in that framework are then particularised to some practically relevant examples in Section 4. Section 5, is devoted to a comparison of the results obtained in Sections 3 and 4 with recent worst case analyzes of IS. Finally, in Section 6, we summarise our main findings and discuss possible venues for future research. Technical proofs and some additional results are provided in the appendices. We conclude this introduction with a brief summary of the notation used throughout the paper.
@@ 2.1 Model
We refer to a pair $\mathcal{M}^{d_{x}}=(\pi_{0},g^{d_{x}})$ as a model for Bayesian inference. Throughout this work, we consider a family of models $\{\mathcal{M}^{d_{x}}\}_{d_{x}\in\mathbb{N}}$, indexed by the dimension $d_{x}$ of the state variable $X$. Intuitively, the models in the family $\{\mathcal{M}^{d_{x}}\}_{d_{x}\in\mathbb{N}}$ may be expected to be “of the same type”, i.e., similar in structure and with differences due mainly to the increasing dimension $d_{x}$. An explicit example (of a family of linear and Gaussian models) is given in Section 4.1. Nevertheless, the conditions required by the main theorems in Sections 3 and 4 actually allow for families where the individual models can be rather different. The dimension of the observation space, $d_{y}$, is assumed to be fixed unless stated otherwise.
@@ 2.2 Importance sampler
For clarity, we carry out our analysis for a simpler version of this general algorithm where the prior is used as the proposal, i.e. $\nu=\pi_{0}$. The resulting scheme is displayed as Algorithm 2. Hereafter, we refer to this specific procedure as standard IS.
@@ 3 Random observations
In this section, we adopt the viewpoint that $Y$ itself is a r.v. generated by the model ${\mathcal{M}}^{d_{x}}$ instead of using it as given (fixed) data. With this perspective, the likelihood function, the normalization constant and the posterior measure also become random objects. In particular, the quantity $\big|\pi_{Y}(f)-\pi_{Y}^{N}(f)\big|$ becomes a $Y$-dependent random variable. The subscript Y emphasizes the randomness of the observation, the same notation is used hereafter for other observation-dependent objects. To derive an upper bound on this random error, we take the expectation w.r.t. the law of $Y$ in order to obtain an averaged error. This leads us naturally to examine how integration over the random observation is performed. The central tool for this development is the Bochner integrability [15] of a specially constructed function, which acts as a link between the fixed and random observation formulations.
@@ 3.1 Random posterior measure
We aim to derive an upper bound for the approximation error in (5), averaged over the observations, where the expectation is taken w.r.t. the joint law of the random samples $x^{1},\ldots,x^{N}$ and the observation $Y$. Much of the analysis to be elaborated in this section depends on two objects that we introduce next:
@@
For any test function $\varphi\in L^{2}(\pi_{0})$, the posterior expectation conditioned on the random observation $Y$ is given by
$\pi_{Y}(\varphi)=\int_{\mathcal{X}}\varphi(x)\,\pi_{Y}({\sf d}x)=\mathbb{E}\!\left[\varphi(X)\mid\mathcal{G}\right],$
@@ 3.3 Approximation errors
    The link function $\ell^{\,d_{x}}$ is Bochner square-integrable, i.e.
    $\ell^{\,d_{x}}\in L^{2}(\mathcal{Y};L^{2}(\pi_{0})).$
@@
$N\;\geq\;\left(\frac{P_{n,p}(d_{x})\|f\|_{\infty}}{\varepsilon}\right)^{2}.$
@@
uniformly w.r.t. the state dimension $d_{x}$.
@@ 4 Special cases
In this section, we examine specific scenarios under which the assumptions of Theorem 3.3 and Corollary 3.4 are satisfied, that is, when the inequalities (14) and (15) holds. It is worth emphasizing that the bounds derived here are not intended to be tight. Instead, our goal is to show that, under the theory developed in the Section 3, the error bounds for the importance sampling estimator $\pi_{Y}^{N}$ may scale polynomially, or even remain uniformly bounded, w.r.t. the state dimension $d_{x}$.
Throughout this Section, we focus on the case $p=2$, which subsumes the case $p=1$
@@
Observe that the quantity
$\|\ell_{y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}=\frac{\pi_{0}([g_{y}^{d_{x}}]^{2})}{\pi_{0}(g_{y}^{d_{x}})^{2}}$
is directly related to the $\chi^{2}$-divergence between the posterior and the prior distributions. It is also related to the relative variance of the Monte Carlo estimator of the normalization constant $Z\mathrel{\mathop{\ordinarycolon}}=\pi_{0}(g_{y}^{d_{x}})$ when sampling from the prior (see Appendix B) and has already been found to be relevant in the analysis of IS methods; see, e.g., [1, 2].
@@
$K_{p}^{d_{x}}=\int_{\mathcal{Y}}\|\ell_{y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{p}\,\eta(\mathrm{d}y)=\int_{\mathcal{Y}}\frac{\|g_{y}^{d_{x}}\|_{L^{2}(\pi_{0})}^{p}}{\pi_{0}(g_{y}^{d_{x}})^{p-1}}\,{\sf d}y.$
@@
$K_{2}^{d_{x}}=\int_{{\mathcal{X}}}\left(\int_{\mathcal{Y}}\frac{g_{y}^{d_{x}}(x)^{2}}{\pi_{0}(g_{y}^{d_{x}})}\,\mathrm{d}y\right)\pi_{0}(\mathrm{d}x).$ (16)
@@
$K_{p}^{d_{x}}=\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{p}\right]<\infty.$ (17)
The previous discussion shows that the key condition required for the dimension-dependent error bounds of Theorem 3.3 to hold reduces to the finiteness of the moment in (17). In other words, the error of the IS estimator is governed by the integrability of the link function $\ell_{Y}^{d_{x}}$ as a random element of $L^{2}(\pi_{0})$. In the remainder of this section, we verify this condition for several relevant observation models. The analysis illustrates how the structure of the likelihood determines the dependence of the error bounds on the state dimension $d_{x}$, and shows that polynomial, or even uniform, bounds can arise in common statistical models. We begin with the linear Gaussian model, since in this case all the relevant distributions can be computed analytically, regardless of the dimensionality of the problem.
@@ 4.1 Linear Gaussian model
$Y=A^{d_{x}}X+V,$
where $A^{d_{x}}\in\mathbb{R}^{d_{y}\times d_{x}}$ is a fixed matrix and $V\sim\mathcal{N}(0,R)$ is a zero-mean Gaussian noise vector with positive definite covariance matrix $R\in\mathbb{R}^{d_{y}\times d_{y}}$, independent of $X$ and also independent of its dimension $d_{x}$. As a result, the r.v. $Y$ is Gaussian with marginal law
$Y\sim\mathcal{N}({\sf d}y;\mu_{y}^{d_{x}},\Sigma_{y}^{d_{x}}),$
@@
$\mu_{y}^{d_{x}}=A^{d_{x}}\mu_{x}^{d_{x}},\qquad\Sigma_{y}^{d_{x}}=A^{d_{x}}\Sigma_{x}^{\,d_{x}}[A^{d_{x}}]^{\top}+R.$ (18)
@@
Hereafter we show that, for the linear Gaussian model described above, the Bochner constant $K_{2}^{d_{x}}$ can exhibit either polynomial growth or remain uniformly bounded w.r.t. the state dimension $d_{x}$, depending on the behavior of the model parameters. In particular, we derive sufficient conditions under which $K_{2}^{d_{x}}$ satisfies a polynomial bound in $d_{x}$, as well as stronger conditions ensuring the existence of a finite constant $\mathcal{K}_{2}$, independent of $d_{x}$, such that inequality (15) holds. Consequently, the assumptions of Corollary 3.4 are satisfied.
4.1.1 Analysis of the Bochner constant $K_{2}^{d_{x}}$
Let us note that, since $A^{d_{x}}\Sigma_{x}^{d_{x}}[A^{d_{x}}]^{\top}$ is positive semidefinite and $R$ is positive definite, it follows that $\Sigma_{y}^{d_{x}}$ in (18) is also positive definite. In addition, the likelihood function $g_{y}^{d_{x}}(x)$ corresponds to the Gaussian pdf of $Y$ conditional on $X=x$, namely,
@@ 4.1.1 Analysis of the Bochner constant $K_{2}^{d_{x}}$
$g_{y}^{d_{x}}(x)^{2}=\frac{1}{(2\uppi)^{d_{y}}|R|}\exp\!\left(-\frac{1}{2}(y-A^{d_{x}}x)^{\top}(2R^{-1})(y-A^{d_{x}}x)\right),$ (20)
which is proportional to a Gaussian density with covariance $\tfrac{1}{2}R$. Indeed, we can express
$g_{y}^{d_{x}}(x)^{2}=C_{1}~\mathcal{N}\!\left(y;A^{d_{x}}x,\tfrac{1}{2}R\right)$ (21)
for some constant $C_{1}$. On the other hand,
$\mathcal{N}\!\left(y;A^{d_{x}}x,\tfrac{1}{2}R\right)=(2\uppi)^{-\frac{d_{y}}{2}}|\tfrac{R}{2}|^{-\frac{1}{2}}\exp\!\left(-\frac{1}{2}(y-A^{d_{x}}x)^{\top}(2R^{-1})(y-A^{d_{x}}x)\right).$ (22)
Combining (20)-(22), and simplifying terms, yields the constant
$C_{1}=(4\pi)^{-d_{y}/2}|R|^{-1/2}.$
Consequently, recalling that $\pi_{0}(\mathrm{d}x)={\mathcal{N}}(\mathrm{d}x;\mu_{x}^{d_{x}},\Sigma_{x}^{d_{x}})$, we readily obtain
$\int_{\mathcal{X}}g_{y}^{d_{x}}(x)^{2}\,\pi_{0}(\mathrm{d}x)=\frac{1}{(4\uppi)^{d_{y}/2}|R|^{1/2}}\mathcal{N}(y;\mu_{y}^{d_{x}},\Sigma_{y}^{d_{x}}-\tfrac{1}{2}R).$ (23)
Using (19) and (23), we realise that, in order to satisfy the integrability condition (16), it suffices to find a finite bound for the expectation
$\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]=\frac{1}{(4\uppi)^{d_{y}/2}|R|^{1/2}}\int_{\mathcal{Y}}\frac{\mathcal{N}(y;\mu_{y}^{d_{x}},\Sigma_{y}^{d_{x}}-\tfrac{1}{2}R)}{\mathcal{N}(y;\mu_{y}^{d_{x}},\Sigma_{y}^{d_{x}})}\,{\sf d}y.$ (24)
Let $S_{1}\mathrel{\mathop{\ordinarycolon}}=\Sigma_{y}^{d_{x}}$ and $S_{2}\mathrel{\mathop{\ordinarycolon}}=\Sigma_{y}^{d_{x}}-\tfrac{1}{2}R$. The integrand in (24) can be rewritten as
$\frac{\mathcal{N}(y;\mu_{y}^{d_{x}},S_{2})}{\mathcal{N}(y;\mu_{y}^{d_{x}},S_{1})}=\left(\frac{|S_{1}|}{|S_{2}|}\right)^{1/2}\exp\!\left(-\tfrac{1}{2}(y-\mu_{y}^{d_{x}})^{\top}(S_{2}^{-1}-S_{1}^{-1})(y-\mu_{y}^{d_{x}})\right).$
@@
$S_{1}=\Sigma_{y}^{d_{x}}\succ\Sigma_{y}^{d_{x}}-\tfrac{1}{2}R=S_{2},$
@@
(hence, $S^{-1}$ is positive definite). The integrand in (24) is thus proportional to the density of a Gaussian distribution with mean $\mu_{y}^{d_{x}}$ and covariance $S$. Then, we can write
$\left(\frac{|S_{1}|}{|S_{2}|}\right)^{1/2}\exp\!\left(-\tfrac{1}{2}(y-\mu_{y}^{d_{x}})^{\top}(S^{-1})(y-\mu_{y}^{d_{x}})\right)=C_{2}\,\mathcal{N}(y;\mu_{y}^{d_{x}},S)$ (26)
for some constant $C_{2}$. On the other hand,
$\mathcal{N}(y;\mu_{y}^{d_{x}},S)=\frac{1}{(2\uppi)^{d_{y}/2}\mathinner{\!\left\lvert S\right\rvert}^{1/2}}\exp\!\left(-\tfrac{1}{2}(y-\mu_{y}^{d_{x}})^{\top}(S^{-1})(y-\mu_{y}^{d_{x}})\right),$ (27)
and combining (26) and (27) we obtain
$C_{2}=\left(\frac{(2\uppi)^{d_{y}}\mathinner{\!\left\lvert S\right\rvert}|S_{1}|}{|S_{2}|}\right)^{1/2}.$ (28)
Therefore, the integral in (24) can be computed exactly to yield
$\int_{\mathcal{Y}}\,\frac{\mathcal{N}(y;\mu_{y}^{d_{x}},S_{2})}{\mathcal{N}(y;\mu_{y}^{d_{x}},S_{1})}\,{\sf d}y=C_{2}\int_{\mathcal{Y}}\,\mathcal{N}(y;\mu_{y}^{d_{x}},S)\,{\sf d}y=C_{2}.$ (29)